OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

: V = πr^2h 300 km^3 = 3.14(7 km)^2 h 1.949824516... = h